# Condensing and Extracting Against Online Adversaries

### v2

@@ Abstract
The existence of condensers (in regimes where extraction is impossible) was recently studied in [CGR, FOCS’24]. They proved condensing impossibility results for various values of $g$ and $\ell$, and they showed the existence of condensers matching the impossibility results in the special case when $n$ is extremely large compared to $\ell$ (i.e., the setting of few blocks of large length).
In this work, we make significant progress on proving the existence of condensers with strong parameters in almost all parameter regimes, even when $n$ is a large enough constant and $\ell$ is growing. This almost resolves the question of the existence of condensers for oNOSF sources, except when $n$ is a small constant.
As our next result, we construct the first explicit condensers for oNOSF sources and achieve parameters that match the existential results of [CGR, FOCS’24]. We also obtain a much improved construction for transforming low-entropy oNOSF sources (where the good blocks only have min-entropy, as opposed to being uniform) into uniform oNOSF sources.
@@
Next, we turn to understanding the possibility of extraction from oNOSF sources. For proving lower bounds, we introduce and initiate a systematic study of a new, natural notion of the influence of functions, which we call online influence, and believe is of independent interest. Using tools from Fourier analysis, we establish tight bounds on the total online influence of functions, which imply extraction lower bounds. Lastly, we give explicit extractor constructions for oNOSF sources, using novel connections to leader election protocols, and further constructing the required leader election protocols. These extractor constructions achieve parameters that go beyond standard resilient functions [AL, Combinatorica’93].
@@ 1 Introduction
Randomness is extremely useful in computation with wide-ranging applications in algorithm design, cryptography, distributed computing protocols, machine learning, error-correcting codes, and much more [motwani1995randomized, vadhan_pseudorandomness_2012]. Most of these applications require access to high quality randomness. However in a lot of settings, especially arising in practice, algorithms only have access to low quality source of randomness. This motivates the notion of condensers: functions that transform weak random sources into strong random sources that are of better quality.
[…]
@@
A function $\mathsf{Cond}:\{0,1\}^{n}\to\{0,1\}^{m}$ is a $(k_{in},k_{out},\varepsilon)$-condenser for a family of distributions $\mathcal{X}$ if for all $\mathbf{X}\in\mathcal{X}$ with $\mathbf{X}\sim\{0,1\}^{n}$ and $H_{\infty}(\mathbf{X})\geq k_{in}$, we have that $H_{\infty}^{\varepsilon}(\mathbf{X})\geq k_{out}$.
@@
The task of the condenser is to make the output entropy rate as high as possible compared to the input entropy rate, or, in other words, to make the output distribution more “condensed”. Related to this, it is also desirable to have as small entropy gap as possible. Notice that if the entropy gap is 00, the output distribution is $\varepsilon$-close to the uniform distribution. Such condensers with entropy gap 00 are known as randomness extractors—a topic that has been extensively studied in theoretical computer science.
When $\mathcal{X}$ is the family of all distributions, it is folklore that no non-trivial condensing is possible. [^1] Assuming $m\leq n$ (wlog this holds since $\left\lvert\mathsf{Cond}(\{0,1\}^{n})\right\rvert\leq 2^{n}$), $m-k_{out}\geq(n-k_{in})-\log(1/(1-\varepsilon))$ and hence the output entropy rate cannot be more than the input entropy rate without incurring extremely large error ($>0.999$). So, we additionally assume that $\mathcal{X}$ is a structured family of sources. [^2] A different route, that has been widely studied, is to assume access to a short independent seed. In this work, we will limit ourselves to the deterministic or seedless setting. Since extractors are the highest quality condensers, a significant amount of work has focused on constructing extractors for interesting family of sources, such as: sources generated by small circuits, two independent sources, algebraically generated sources, sources generated by small space sources, and many more [tv00, chattopadhyay_explicit_2019, dgw09extractpolynomials, kamp_deterministic_2007].
However, for many natural family of sources, one can provably show that no extractor can exist. In such situations, one can still hope to show that high quality condensers exist. We note that condensers (and sources with high min-entropy rate) are very useful: the condensed distribution can be used to efficiently simulate randomized algorithms with small overhead, perform one-shot simulations for randomized protocols, cryptography and interactive proofs, and much more. [DPW14key] showed these condensers are equivalent to ‘unpredictability extractors’ that can simulate cryptographic protocols against biased distinguishers. For details on these applications and more, see [aggarwal_how_2020, doron_almost_2023, CGR_seedless_condensers].
In this work, we focus on one natural family of sources where it is known that extraction is impossible (for many interesting parameter regimes). The family we consider are known as online non-oblivious symbol fixing sources (oNOSF sources ). [^3] These sources are in contrast to non-oblivious symbol fixing (NOSF) sources where bad blocks can be arbitrary functions of all the good blocks. NOSF sources were introduced in [chor_bit_1985] with applications in leakage-resilient cryptography, and have been well-studied. Formally:
@@
A ${\left(g,\ell,n,k\right)}$-oNOSF source $\mathbf{X}=(\mathbf{X}_{1},\dots,\mathbf{X}_{\ell})$ is such that each block $\mathbf{X}_{i}$ is over $\{0,1\}^{n}$, $g$ of the blocks are are independent sources with min-entropy $k$ (“good blocks”), and each “bad block” is an arbitrary function of the blocks with an index smaller than it. When $k=n$, we will call such sources uniform ${\left(g,\ell,n\right)}$-oNOSF sources.
These sources are inspired by real-time randomness generation settings such as in blockchains. There, each subsequent block is random or controlled by an adversary. Since these sources are generated in real time, a bad block can only be a function of the blocks that have appeared so far, and it is reasonable to assume that the good blocks contain entropy and are independent. Further, there are natural cryptographic settings, such as creating a Common Reference String, that are widely used in various cryptographic protocols where oNOSF source sources naturally arise (see [aggarwal_how_2020] for a discussion).
@@
We almost fully resolve the question of the existence of condensers for oNOSF sources by showing that good condensers, meeting the impossibility results of [CGR_seedless_condensers], exist when the number of blocks $\ell$ is growing and $n$ is a large constant. We also construct explicit condensers for oNOSF sources matching the results of [CGR_seedless_condensers] and obtain an improved construction for transforming low-entropy oNOSF sources into uniform oNOSF sources . Moreover, we find new applications of our results on condensers to collective coin flipping and collective sampling, and use these connections to provide simple protocols for these problems. In the context of extractors for oNOSF sources , we introduce the new, natural notion of online influence for Boolean functions and show extraction lower bounds for oNOSF sources by establishing tight bounds on the total online influence of functions. Lastly, using our novel connections to leader election protocols, we construct explicit extractors for uniform oNOSF sources by explicitly constructing the required leader election protocols, the results of which are summarized in Tables 1 and 2.
@@
The remainder of our introduction is structured as follows. We give an overview of previous work in Section 1.1 before presenting our main existential and explicit condenser results in Section 1.2. In Section 1.3, we present our results on the limits of extraction from oNOSF sources . In Section 1.4, we show how our results on condensers have implications for collective coin flipping and sampling protocols.
@@ 1.1 Previous Work
The study of extractors for oNOSF sources was initiated by [aggarwal_how_2020]. [^4] In [aggarwal_how_2020], these sources were called SHELA (Somewhere Honest Entropic Look Ahead) sources. Their results include the following:
@@
  • An explicit transformation from ${\left(g,\ell,n,0.9n\right)}$-oNOSF source into a source over ($(\{0,1\}^{O(n)})^{\ell-1}$ where $g-1$ of the blocks are uniform and independent.
  • An explicit transformation from ${\left(g,\ell,n,0.1n\right)}$-oNOSF source into a source over ($(\{0,1\}^{O(n)})^{100\ell}$ where $g-1$ of the blocks are uniform and independent.
Even though the output entropy rate is only slightly more than the input-entropy rate in the second result and smaller in the third result, the fact that a lot of the blocks are truly uniform is very useful, and they find interesting cryptographic applications of these somewhere-extractors.
Before our work, the best known extractors for oNOSF sources could be obtained by using resilient functions or equivalently, extractors for NOSF sources (non-online version of oNOSF sources ) constructed by [ajtai_influence_1993, chattopadhyay_explicit_2019, meka_explicit_2017, IMV23resilient, IV24resilient] ; these require $g\geq\ell-\frac{\ell}{(\log\ell)^{2}}$.
@@
oNOSF sources were further studied by [CGR_seedless_condensers], where they obtained the following results regarding condensers:
@@
  • When $n\geq 2^{\omega(\ell)}$, there exists condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m=O(n\cdot\ell/g)}$ such that for any uniform ${\left(g,\ell,n\right)}$-oNOSF source $\mathbf{X}$, $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq\frac{1}{\left\lfloor\ell/g\right\rfloor}\cdot m-O(\log(n/\varepsilon))$. Their result is not explicit.
[…]
We also mention a related family of sources, namely adversarial Chor-Goldreich sources. Uniform oNOSF sources can be seen as a special case of adversarial Chor-Goldreich sources where the good blocks are uniform. Constructing condensers where the output entropy rate is $g/\ell$ for adversarial Chor-Goldreich sources is already a challenging task, although such condensers in various parameter regimes have been recently constructed [doron_almost_2023, GLZ_cg_condenser]. The paper of [DMOZ24condenseunpredictability] recently constructed condensers for a related more general model.
1.2 New Condenser Constructions
Previous works only showed the existence of condensers for oNOSF sources when $n\geq 2^{\omega(\ell)}$. We vastly improve on this result in two ways. First, we show that for almost all values of $n,\ell$, even when $n$ is a small constant, excellent condensers exist. Second, we provide explicit condensers for oNOSF sources when $n\geq 2^{\omega(\ell)}$. We also obtain much better transformation from low-entropy oNOSF sources to uniform oNOSF sources that work even when $k\ll\ell$. These results show condensers always exist, except when $n$ is a very small constant (such as $n=1$). To further our understanding of this case, we initiate the study of online influence of Boolean functions, a natural generalization of influence that captures the one-sided nature of our online adversary. We also discover surprising connections between condensers for oNOSF sources and protocols for natural problems in distributed computing, such as collective coin flipping and collective sampling. We now discuss our result in details below.
1.2.1 Existential Condensers
@@
For all $\ell,\varepsilon$ where $\ell\geq O(\log(1/\varepsilon))$, and $n=10^{4}$, there exists a condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any uniform ${\left(0.51\ell,\ell,n\right)}$-oNOSF source $\mathbf{X}$, we have $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq 0.99m$ where $m=\Omega(\ell+\log(1/\varepsilon))$. Furthermore, when $n=\omega(1)$, the output entropy rate becomes $1-o(1)$.
This is tight since [CGR_seedless_condensers] showed it is impossible to condense uniform ${\left(0.5\ell,\ell,n\right)}$-oNOSF sources beyond output entropy rate $0.5$.
Theorem 2.
For all $\ell,n,\varepsilon$ where $n=\operatorname{poly}(\log(\ell/\varepsilon)),k=O(\log(\ell/\varepsilon))$, there exists a condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any ${\left(0.51\ell,\ell,n,k\right)}$-oNOSF source $\mathbf{X}$, we have $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-O(m/\log(m))-O(\log(1/\varepsilon))$ where $m=\Omega(k)$.
We can also extend our result to condense from uniform ${\left(g,\ell,n\right)}$-oNOSF sources for all $g,\ell$ and constant $n$ where the output entropy rate is $0.0011/\left\lfloor\ell/g\right\rfloor-$ 0.001. This is tight since [CGR_seedless_condensers] showed it is impossible to condense such sources beyond output entropy rate $1/\left\lfloor\ell/g\right\rfloor$.
Previously, [CGR_seedless_condensers] showed how to existentially condense from uniform ${\left(g,\ell,n\right)}$-oNOSF sources when $g\geq 0.51\ell$, provided $n\geq 2^{\omega(\ell)}$. As $n$ gets smaller, condensing becomes harder since a uniform ${\left(g,\ell,n\right)}$-oNOSF source is also a uniform ${\left(g\cdot n/1000,\ell\cdot n/1000,1000\right)}$-oNOSF source. Hence, we greatly improve the parameters while using different and much simpler techniques.
1.2.2 Explicit Condensers
We construct the first explicit condensers for oNOSF sources . Our explicit condenser construction achieves the same parameters as the existential condenser construction of [CGR_seedless_condensers]. We show how to explicitly condense from uniform ${\left(g,\ell,n\right)}$-oNOSF sources when $n\geq 2^{\omega(\ell)}$ and $g\geq 0.5\ell+1$. We state the results for constant $\ell$ since that is cleaner:
Theorem 3 (Informal version of Theorem 5.1).
For all $n,\varepsilon$ and constant $\ell$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any uniform ${\left(0.5\ell+1,\ell\right)}$-oNOSF source $\mathbf{X}$, we have $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-O(\log(m/\varepsilon))$, where $m=\Omega(n)$.
Just like earlier, since condensing when $g=0.5\ell$ is impossible, this result is also tight. Using our new results regarding transforming oNOSF sources to uniform oNOSF sources , we also obtain explicit condensers for ${\left(0.51\ell,\ell,n,k\right)}$-oNOSF sources for the same parameter regime:
@@ 1.2.2 Explicit Condensers
For all $\ell,n,\varepsilon$ with constant $\ell$ and $n\geq O(\log(1/\varepsilon))$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any ${\left(0.5\ell+2,\ell,n,\operatorname{poly}(\log n)\right)}$-oNOSF source $\mathbf{X}$, we have $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-O(\log(m/\varepsilon))$ where $m=\operatorname{poly}(\log n)$.
Similar to earlier, we can also extend our result to explicitly condense from uniform ${\left(g,\ell,n\right)}$-oNOSF sources in the same parameter regime so that the output entropy rate is $1/\left\lfloor\ell/g\right\rfloor-o(1)$. Just like earlier, this is tight as well.
Previously, [CGR_seedless_condensers] showed how to existentially condense from uniform ${\left(g,\ell,n\right)}$-oNOSF sources in this parameter regime. However, they relied on the existence of a very strong pseudorandom object: “output-light” low-error two-source extractors. Such extractors, even without the output-lightness requirement, are extremely hard to construct and it is a major open problem to obtain such extractors. We are able to make this condenser explicit by building up on their ideas, making interesting observations regarding oNOSF sources , and stitching them together so that the base pseudorandom object we rely on are seeded extractors that we know how to explicitly construct with near optimal parameters.

### v3

@@ Abstract
The existence of condensers (in regimes where extraction is impossible) was recently studied in [CGR, FOCS’24]. They proved condensing impossibility results for various values of $g$ and $\ell$, and they showed the existence of condensers matching the impossibility results in the special case when $n$ is exponential in $\ell$ (i.e., the setting of few blocks of large length).
In this work, not only do we construct the first explicit condensers matching the existential results of [CGR, FOCS’24], but we make a doubly exponential improvement by handling the case when $n$ is only polylogarithmic in $\ell$. We also obtain a much improved explicit construction for transforming low-entropy oNOSF sources (where the good blocks only have min-entropy, as opposed to being uniform) into uniform oNOSF sources.
As our next result, we essentially resolve the question of the existence of condensers for oNOSF sources by showing the existence of condensers in almost all parameter regimes, even when $n$ is a large enough constant and $\ell$ is growing.
@@
Next, we turn to understanding the possibility of extraction from oNOSF sources. For proving lower bounds, we introduce and initiate a systematic study of a new, natural notion of the influence of functions, which we call online influence, and establish tight bounds on the total online influence of functions, which imply extraction lower bounds. Lastly, we give explicit extractor constructions for oNOSF sources using novel connections to leader election protocols, and we further construct the required leader election protocols. These extractor constructions achieve parameters that go beyond the standard resilient functions of [AL, Combinatorica’93].
Contents
@@ 1 Introduction
Randomness is extremely useful in computation with wide-ranging applications in algorithm design, cryptography, distributed computing protocols, machine learning, error-correcting codes, and much more [MR95, Vad12]. Most of these applications require access to high quality randomness. However in a lot of settings, especially arising in practice, algorithms only have access to low quality source of randomness. This motivates the notion of condensers: functions that transform weak random sources into strong random sources that are of better quality.
The standard way of measuring the amount of randomness is using min-entropy. Formally, for a source (distribution) $\mathbf{X}$ with support $\Omega$, define its min-entropy as $H_{\infty}(\mathbf{X})=\min_{x\in\Omega}\log_{2}(1/\Pr[\mathbf{X}=x])$. We will also need the notion of smooth min-entropy, which measures how close a distribution is to having high entropy. Formally, for a source $\mathbf{X}$, its smooth min-entropy with parameter $\varepsilon$ is defined as $H_{\infty}^{\varepsilon}(\mathbf{X})=\max_{\mathbf{Y}:\left\lvert\mathbf{X}-\mathbf{Y}\right\rvert\leq\varepsilon}\{H_{\infty}(\mathbf{Y})\}$, where $\left\lvert\cdot\right\rvert$ denotes the statistical distance (Definition 4.1).
@@
A function $\mathsf{Cond}:\{0,1\}^{n}\to\{0,1\}^{m}$ is a $(k_{in},k_{out},\varepsilon)$-condenser for a family of distributions $\mathcal{X}$ if for all $\mathbf{X}\in\mathcal{X}$ with $\mathbf{X}\sim\{0,1\}^{n}$ and $H_{\infty}(\mathsf{Cond}(\mathbf{X}))\geq k_{in}$, we have that $H_{\infty}^{\varepsilon}(\mathbf{X})\geq k_{out}$.
@@
The task of the condenser is to make the output entropy rate as high as possible compared to the input entropy rate; i.e., to make the output distribution more “condensed”. Related to this, it is also desirable to have as small entropy gap as possible. Condensers with entropy gap 0 are known as randomness extractors and have been extensively studied in theoretical computer science.
When $\mathcal{X}$ is the family of all distributions, it is folklore that no non-trivial condensing is possible. [^1] Assuming $m\leq n$ (wlog this holds since $\left\lvert\mathsf{Cond}(\{0,1\}^{n})\right\rvert\leq 2^{n}$), $m-k_{out}\geq(n-k_{in})-\log(1/(1-\varepsilon))$ and hence the output entropy rate cannot be more than the input entropy rate without incurring extremely large error (>0.999>0.999). So, we additionally assume that $\mathcal{X}$ is a structured family of sources. [^2] A different route, that has been widely studied, is to assume access to a short independent seed. In this work, we will limit ourselves to the deterministic or seedless setting. Since extractors are the highest quality condensers, a significant amount of work has focused on constructing extractors for many interesting families of sources [TV00, CZ19, DGW09, KZ07]. However, for many natural family of sources, one can provably show that no extractor can exist.
In this work, we focus on one natural family of sources where it is known that extraction is impossible (for many interesting parameter regimes): online non-oblivious symbol fixing sources (oNOSF sources ). [^3] These sources are in contrast to non-oblivious symbol fixing (NOSF) sources where bad blocks can be arbitrary functions of all the good blocks. NOSF sources were introduced in [CGHFRS85] with applications in leakage-resilient cryptography, and have been well-studied. Formally:
@@
A $\left(g,\ell,n,k\right)$}-oNOSF source $\mathbf{X}=(\mathbf{X}_{1},\dots,\mathbf{X}_{\ell})$ is such that each block $\mathbf{X}_{i}$ is over $\{0,1\}^{n}$, $g$ of the blocks are independent sources with min-entropy $k$ (“good blocks”), and each “bad block” is an arbitrary function of the blocks with an index smaller than it. When $k=n$, we will call such sources uniform $\left(g,\ell,n\right)$}-oNOSF sources.
@@
The previous work of [CGR24] gave a condenser impossibility result for oNOSF sources and showed the existence of a condenser matching that result as long as the block length $n$ was exponential in the number of blocks $\ell$. We construct explicit condensers for oNOSF sources that match the results of [CGR24]; in fact, we only require $n$ to be polylogarithmic in $\ell$, providing a doubly exponential improvement over the (existential) result of [CGR24]. Next, we essentially resolve the existence question for oNOSF source condensers by showing that good condensers exist even when $n$ is a (large) constant and $\ell$ grows. To go with these results, we obtain an improved construction for transforming low-entropy oNOSF sources into uniform oNOSF sources . Moreover, we find new applications of our results on condensers to collective coin flipping and collective sampling, and use these connections to provide simple protocols for these problems. In the complementary direction, we construct explicit extractors for uniform oNOSF sources by explicitly constructing the required leader election protocols, the results of which are summarized in Tables 1 and 2. Also in the context of extractors for oNOSF sources , we introduce the new, natural notion of online influence for Boolean functions and show extraction lower bounds for oNOSF sources by establishing tight bounds on the total online influence of functions.
On the Utility of Condensing for oNOSF sources
We note that condensers (and sources with high min-entropy rate) are very useful: the condensed distribution can be used to efficiently simulate randomized algorithms with small overhead, perform one-shot simulations for randomized protocols, cryptography, interactive proofs, and much more. For details on these applications and more, see [AORSV20, DMOZ23, CGR24, DPW14].
Practical applications to blockchains and cryptography
oNOSF sources are inspired by real-time randomness generation settings such as in blockchains where the adversary has some probability of corrupting a block. Moreover, it is known that non-corrupted blocks have some amount of min-entropy [BCG15]. In fact, several works have attempted to use Bitcoin or Ethereum as a source of randomness in cryptographic protocols [BCG15, BGZ16, PW18, BGB17]. However, the authors of [BGZ16] showed that even when the adversary has a small, constant probability of corrupting a block, randomness extraction is impossible from Bitcoin. [^4] This mirrors our extraction impossibility result for oNOSF sources in Section 10 Our results show that in this setting, it is still possible to get a condensed source with a high min-entropy rate. It is known that such sources are still useful for cryptographic protocols, such as hedged public-key enryption [BBNRSSY09]. Further, there are natural cryptographic settings, such as creating a Common Reference String, that are widely used in various cryptographic protocols where oNOSF sources naturally arise [AORSV20].
Practical applications to fault-tolerant distributed computing
One common scenario in distributed computing is that of many agents (e.g., servers in a network) attempting to collectively take a decision using several rounds of communication over a common broadcast channel in the presence of computationally unbounded adversarial agents, which render cryptographic primitives ineffective. Protocols for collective coin flipping, leader election, and collective sampling are prime examples of this scenario that have been intensively studied ([BL89, GGL91, Dod06, AN93, Fei99] and many more). In Section 3, we explain how condensing or extracting from oNOSF sources can be viewed as a variant of these protocols. As a consequence, our new results on condensers provide a new protocol for collective sampling and impossibility results for these protocols can be translated into lower bounds against extractors and condensers for oNOSF sources .
@@
The remainder of our introduction is structured as follows. We give an overview of previous work in Section 1.1 before presenting our main existential and explicit condenser results in Section 1.2. In Section 1.3, we present our results on the limits of extraction from oNOSF sources . Later on, in Section 3, we explain how our results on condensers have implications for collective coin flipping and sampling protocols.
@@ 1.1 Previous Work
The study of extractors for oNOSF sources was initiated by [AORSV20]. [^5] In [AORSV20], these sources were called SHELA (Somewhere Honest Entropic Look Ahead) sources. Their results include the following:
@@
  • An explicit transformation from $\left(g,\ell,n,0.9n\right)$}-oNOSF source into a source over $(\{0,1\}^{O(n)})^{\ell-1}$ where g1g-1 of the blocks are uniform and independent.
  • An explicit transformation from $\left(g,\ell,n,0.1n\right)$}-oNOSF source into a source over $(\{0,1\}^{O(n)})^{100\ell}$ where g1g-1 of the blocks are uniform and independent.
Even though the output entropy rate is only slightly more than the input-entropy rate in the second result and smaller in the third result, the fact that a lot of the blocks are truly uniform is very useful, and they find interesting cryptographic applications of these “somewhere-extractors”.
Before our work, the best known extractors for oNOSF sources could be obtained by using resilient functions or equivalently, extractors for NOSF sources (non-online version of oNOSF sources ) constructed by [AL93, CZ19, Mek17, IMV23, IV24] ; these require $g\geq\ell-\frac{\ell}{(\log\ell)^{2}}$.
@@
oNOSF sources were further studied by [CGR24], where they obtained the following results regarding condensers:
@@
  • When $n\geq 2^{\omega(\ell)}$ and $g>0.5\ell$ , there exists condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m=O(n\cdot\ell/g)}$ such that for any uniform $\left(g,\ell,n\right)$}-oNOSF source $\mathbf{X}$, $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-O(\log(n/\varepsilon))$. [^6] They get a tradeoff for $g\leq 0.5\ell$ as well Their result is not explicit.
  • It is impossible to condense from $\textrm{uniform }{\left(0.5\ell,\ell,n\right)}\textrm{-oNOSF sources}$ with output entropy rate more than 0.50.5.777They get impossibility for other smaller $g$ as well
We also mention a related family of sources, namely adversarial Chor-Goldreich sources. Uniform oNOSF sources can be seen as a special case of adversarial Chor-Goldreich sources where the good blocks are uniform. Constructing condensers where the output entropy rate is $g/\ell$ for adversarial Chor-Goldreich sources is already a challenging task, although such condensers in various parameter regimes have been recently constructed [DMOZ23, GLZ24]. The paper of [DMOZ25] recently constructed condensers for a related, more general model.
1.2 This Work: New Condenser Constructions
Previous works only showed the existence of condensers for oNOSF sources when $n\geq 2^{\omega(\ell)}$. We vastly improve on this result in two ways. First, we construct explicit condensers that work even when $n\geq\operatorname{polylog}(\ell)$ and provide an explicit transformation from low-entropy oNOSF sources to uniform oNOSF sources that works even when the min-entropy of a block $k$ is only $\operatorname{polylog}(n)$. Second, we show that condensers for oNOSF sources exist when $n$ is just a large constant, only leaving open the question of the existence of such condensers for when $n$ is a very small constant (e.g., $n=1$). We also discover surprising connections between condensers for oNOSF sources and protocols for natural problems in distributed computing, such as collective coin flipping and collective sampling. Lastly, we initiate the study of online influence of Boolean functions, a natural generalization of influence that captures the one-sided nature of our online adversary to help us analyze the setting of $n=1$. We now discuss our results in detail below.
1.2.1 Explicit Condensers
We construct the first explicit condensers for oNOSF sources . Our ultimate result is founded on a baseline construction that itself is an explicit condenser for oNOSF sources that matches the existential results of [CGR24] and works for any block length $n=2^{\Omega(\ell)}\log(1/\varepsilon)$ as long as at least 51% of blocks are good.
Theorem 1 (Theorem 2.4 restated).
For all $\varepsilon>0$ and $n,\ell\in\mathbb{N}$ where $n\geq 2^{\Omega(\ell)}\log(1/\varepsilon)$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any $(g=0.51\ell,\ell,n)$-oNOSF source $\mathbf{X}$, we have that $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-2^{O(\ell)}\log(1/\varepsilon)$ where $m=0.0001\ell n$.
Surprisingly, we are able to improve upon this baseline to obtain explicit condensers that work for oNOSF sources where the block length $n$ is only at least $\operatorname{poly}(\log(\ell)/\varepsilon)$.
@@ 1.2.1 Explicit Condensers
For all $\varepsilon>0$ and $n,\ell\in\mathbb{N}$ where $n\geq\operatorname{poly}(\log(\ell)/\varepsilon)$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any $(g=0.51\ell,\ell,n)$-oNOSF source $\mathbf{X}$, we have that $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-\operatorname{poly}(\log(\ell)/\varepsilon)\cdot\log(n)$ where $m=0.001\ell n-O(\ell\log(\ell)\log(1/\varepsilon))$.
Since condensing when $g=0.5\ell$ is impossible, both results are tight. We note that neither result completely subsumes the other. Our baseline construction in Theorem 1 has an exponential dependence of $n$ on $\ell$ instead of the polylogarithmic dependence achieved in Theorem 2; however, the latter result requires the dependence $n\geq\operatorname{poly}(1/\varepsilon)$ compared to $n\geq\log(1/\varepsilon)$ for the baseline construction.
Using our new results regarding transforming oNOSF sources to uniform oNOSF sources , we also obtain explicit condensers for $\left(0.51\ell,\ell,n,k\right)$}-oNOSF sources for the same parameter regime:
@@
For all $n,\ell,k\in\mathbb{N}$ where $n\geq\operatorname{poly}(\ell)$ and $k\geq\operatorname{polylog}(n)$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any $(g=0.51\ell,\ell,n,k)$-oNOSF source $\mathbf{X}$, we have that $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-\operatorname{polylog}(\ell)\cdot\log(n)$ where $m=0.001\ell n-O(\ell\log(\ell)\log\log(\ell))$ and $\varepsilon=\operatorname{poly}(1/\log(\ell))$.
We can also extend our result to explicitly condense from uniform $\left(g,\ell,n\right)$}-oNOSF sources in the same parameter regime so that the output entropy rate is $1/\left\lfloor\ell/g\right\rfloor-o(1)$, which is tight according to the impossibility result of [CGR24].
Previously, [CGR24] showed how to existentially condense from uniform $\left(g,\ell,n\right)$}-oNOSF sources when $n=2^{\Omega(\ell)}$. However, they relied on the existence of a very strong pseudorandom object: “output-light” low-error two-source extractors. Such extractors, even without the output-lightness requirement, are extremely hard to construct and it is a major open problem to obtain such extractors. We are able to construct explicit condensers by creating new tools that allow us to use an oNOSF source to sample indices within an oNOSF source , and stitching them together so that the base pseudorandom object we rely on are seeded extractors that we know how to explicitly construct with near optimal parameters.